# A BERTology View of LLM Orchestrations: Token- and Layer-Selective Probes for Efficient Single-Pass Classification

### v1

@@ A.1 Search Ranges
Attention heads $\{4,8,16\}$
Downcast factor $\{4,8,16,32,64\}$
Table 6: Hyperparameter search ranges. The downcast factor determines the inner attention dimension as $d_{\text{inner}}=d/\text{factor}$, where $d=3072$ for Llama-3.2-3B.
@@ A.2 Experimental Setup
All experiments were conducted on a single NVIDIA RTX 3090 GPU (24GB VRAM) with 96GB system RAM. To reduce peak GPU memory during training, we pre-extracted and cached hidden states from the frozen LLM before training probe classifiers, allowing larger batch sizes under VRAM constraints.
@@ Appendix B Ablation Study on Attention Downcasting
We evaluate the effect of dimension downcasting on the multi-head self-attention aggregation mechanism using the ToxicChat dataset. The results are shown on Table 7. The downcasting factor determines the inner dimension $d_{\text{inner}}=d/\text{factor}$ for the QKV projections, where $d$ is the hidden dimension (d=3,072d=3{,}072 for Llama-3.2-3B).

### v2

@@ A.1 Search Ranges
Attention heads $\{4,8,16\}^{\dagger}$
Downcast factor $\{4,8,16,32,64\}^{\ddagger}$
[…]
@@ A.2 Experimental Setup
All probe training was conducted on a single NVIDIA RTX 3090 GPU (24GB VRAM) with 96GB system RAM. For the additional backbones (GPT-OSS-20B and Qwen3-30B-A3B), hidden states were pre-extracted on an NVIDIA A100 80GB GPU and transferred for local probe training. To reduce peak GPU memory during training, we pre-extracted and cached hidden states from the frozen LLM before training probe classifiers, allowing larger batch sizes under VRAM constraints.
@@ Appendix B Ablation Study on Attention Downcasting
We evaluate the effect of dimension downcasting on the multi-head self-attention aggregation mechanism using the ToxicChat and SST2 datasets. The results are shown in Table 7. The downcasting factor determines the inner dimension $d_{\text{inner}}=d/\text{factor}$ for the QKV projections, where $d$ is the hidden dimension (d=3,072d=3{,}072 for Llama-3.2-3B).